using f(x)=3x-1 and g(x)=1-x^2
solve 1.
f(g(0)): g(0) is 1-0^2 = 1
Plugging the 1 in f(x) you get f(1)= 3x1 - 1 = 2
So f(g(0))= 2
Solve 2:
Similarly f(2)= 3x2 -1 = 5
And plugging that value 5 in g(5) you get 1- 5^2 = -24
1.f(g(0))
2.g(f(2))

Answer by muthbab@gmail.com(5)

using f(x)=3x-1 and g(x)=1-x^2
solve 1.
f(g(0)): g(0) is 1-0^2 = 1
Plugging the 1 in f(x) you get f(1)= 3x1 - 1 = 2
So f(g(0))= 2
Solve 2:
Similarly f(2)= 3x2 -1 = 5
And plugging that value 5 in g(5) you get 1- 5^2 = -24
1.f(g(0))
2.g(f(2))